# Pullback of arithmetic theta series and its modularity for unitary Shimura curves

### v1

@@ 6.2. Arithmetic divisors
By Theorem 5.5, we have
$\varphi_{\Lambda}^{*}(\mathcal{Z}^{\mathrm{tot},\diamond}(m,v))=\sum_{m_{1}+m_{2}=m}r_{\Lambda}(m_{2})\cdot\mathcal{Z}^{\mathrm{tot}}(m_{1},v).$
It remains to check the Green functions of the divisors from the two sides agree. Since both Green functions are analytic, it suffices to check this on the open Shimura variety $\mathcal{S}(\mathbb{C})$. It is easy to see by simply unfolding (the Green functions) that it suffices to prove the following local identity for $z\in\mathcal{D}$
(6.13) $\mathrm{Gr}^{\diamond}(x,v)(z)=\mathrm{Gr}(x_{1},v)(z)$
for any $x=x_{1}+x_{2}$ with $x_{1}\in V_{\mathbb{R}}$ and $x_{2}\in\Lambda_{\mathbb{R}}$, except that some modification is needed for $x_{1}=0$. When $x_{1}\neq 0$, (6.13) is obvious as $R(x,z)=R(x_{1},z)$ for $z\in\mathcal{D}$. This finishes the case with $m\neq 0$ and $x_{1}\neq 0$. The case of $m=0$ follows basically from definition.
So we are left with the case $m>0$ and $x\in\Lambda_{\mathbb{R}}$ with $(x,x)=m$. Viewing $\mathcal{D}^{\diamond}(x)$ as a line bundle over $\mathcal{D}^{\diamond}$ with the metric given by $\mathrm{Gr}^{\diamond}(x,v)$, we have a line bundle $\mathcal{R}=\mathcal{D}^{\diamond}(x)\otimes\Omega$ with product metric. By Lemma 5.4, $\mathcal{R}|_{\mathcal{D}}$ is a trivial line bundle. We need to verify that the product metric on $\mathcal{R}$ is constant $1$ when restricting to $\mathcal{D}$.
For $0\neq x\in\Lambda_{\mathbb{R}}$, we have $(x,x)>0$ and we can define a section $s_{x}$ of $\Omega^{-1}$ via
(6.14) $s_{x}(a)=(x,a),\quad\text{where $a$ is a section of $\Omega|_{U}$}.$
Over a point $z\in\mathcal{D}^{\diamond}$, we have
$s_{x,z}(a)=(x,a)=(\mathrm{pr}_{z}(x),a),\quad\text{ for }a\in z.$
@@
(6.15) $-\log\|s_{x,z}\|^{2}=-\gamma-\log(4\pi vR(x,z)).$
On the other hand, the canonical section ${\bf 1}$ of the line bundle $\mathcal{D}^{\diamond}(x)$ has self inner product
$-\log\|{\bf 1}_{z}\|^{2}=\mathrm{Gr}^{\diamond}(x,v)=\beta_{1}(4\pi vR(x,z))=-\gamma-\log(4\pi vR(x,z))+f_{x}(z)$
with a smooth function
$f_{x}(z)=-\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}\frac{(4\pi v\cdot R(x,z))^{k}}{k!}.$
So the section $r={\bf 1}\otimes s_{x}^{-1}$ on $\mathcal{R}$ has the following property
$-\log\|r_{z}\|^{2}=\mathrm{Gr}^{\diamond}(x,z)+\log\|s_{x,z}\|^{2}=f_{x}(z),$
which vanishes when $z\in\mathcal{D}\subset\mathcal{D}^{\diamond}(x)$. That is $\|r_{z}\|=1$ for $z\in\mathcal{D}$ as expected. This proves (6.12) for Green function part on $\mathcal{S}$. This proves the theorem. ∎
@@ 7.1. Eisenstein Series
[…]
$\lambda:\,S(V_{\mathbb{A}})\rightarrow I(s_{0},\chi),\quad\lambda(\phi)(g)=\omega(g)\phi(0).$
Here $\chi=\chi_{-D}^{\ell}$. Let $\phi_{L}=\mathrm{Char}(\widehat{L})\otimes\phi_{\infty}\in S(V_{\mathbb{A}})$ with $\phi_{\infty}(x)=e^{-\pi(x,x)}$, and let $\Phi=\Phi_{L}=\prod_{p\leq\infty}\Phi_{p}$ be the standard section in $I(s,\chi)$ such that $\Phi_{L}(g,s_{0})=\lambda(\phi_{L})(g)$. Then it is known that $\Phi_{\infty}=\gamma(V_{\infty})\Phi_{\infty}^{\ell}$ is the standard section of weight $\ell$ (up to a scalar multiplication by a local Weil index $\gamma(V_{\infty})$):
@@
Let
(7.1) $E_{L}(\tau,s)=v^{-\ell/2}\sum_{\gamma\in B\backslash\operatorname{SL}_{2}(\mathbb{Q})}\Phi_{L}(\gamma g_{\tau},s),\quad\tau=u+iv\in\mathbb{H}$
be the associated Eisenstein series of weight $k$ (level $D$ and character $\chi$), which depends only on the genus $[[L]]$ of $L$. Here
@@
Let $\Gamma_{\infty}=B\cap\operatorname{SL}_{2}(\mathbb{Z})$, then $B\backslash\operatorname{SL}_{2}(\mathbb{Q})=\Gamma_{\infty}\backslash\operatorname{SL}_{2}(\mathbb{Z})$.
@@
$\Phi_{\infty}^{\ell}(\gamma g_{\tau},s_{0})=v^{\frac{\ell}{2}}j(\gamma,\tau)^{-\ell}.$
@@
    When $p\nmid D$, $\Phi_{p}(\gamma,s_{0})=1$ and $\Phi_{p}$ is right $\operatorname{SL}_{2}(\mathbb{Z}_{p})$-invariant.
@@
    $\Phi_{p}(\gamma,s_{0})=\begin{cases}\chi_{p}(a)&\hbox{if }p|c,\\ \gamma(V_{p})\chi_{p}(c)p^{-\ell/2}&\hbox{if }p\nmid c.\end{cases}$
@@
Notice that similar results hold when $s_{0}$ is replaced by general $s$ although we do not need it here. We have now the following theorem by the above lemmas.
@@
Let the notation be as above. Then for $\ell\geq 4$ and $s_{0}=\ell-1$,
$E_{L}(\tau,s_{0})$ $=\sum_{\gamma\in\Gamma_{\infty}\backslash\operatorname{SL}_{2}(\mathbb{Z})}a(\gamma,\ell)j(\gamma,\tau)^{-\ell}$
$=\sum_{\delta\in\Gamma_{\infty}\backslash\operatorname{SL}_{2}(\mathbb{Z})/\Gamma_{1}(D)}a(\delta,\ell)\sum_{\gamma\in\Gamma_{\infty}\backslash\delta\Gamma_{1}(N)}j(\gamma,\tau)^{-\ell}.$
@@ 7.2. Modularity of arithmetic theta functions with Kudla Green functions
Let the notation be as above. Then $\widehat{\Theta}_{K}(\tau)$ is a (non-holomorphic) modular form of $\Gamma_{0}(D)$ with values in $\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$.
@@
Here $s_{\ell}=\ell-1$. By [${}^{+}$20, Theorem B] and [ES18, Theorem 1.4], the left hand side of (7.3) is a (non-)holomorphic modular form of $\Gamma_{0}(D)$ of weight $2+\ell$ character $\chi_{-D}^{\ell+2}$, we see that $\widehat{\Theta}_{K}(\tau)$ transforms like a modular form of weight $2$ for $\Gamma_{0}(D)$. We still need to prove that $\widehat{\Theta}_{K}(\tau)$ is well-defined everywhere. By Theorem 7.3, we see that
$E_{L_{0}^{\ell}}(\tau,s_{\ell})=\sum_{\gamma\in\Gamma_{\infty}\backslash\mathrm{SL}_{2}(\mathbb{Z})}(\prod_{p|D}b_{p}(\gamma,\ell)\cdot j(\gamma,\ell)^{-1})^{\ell},$
@@
Now, the modularity for $\widehat{\Theta}_{B}(\tau)$ should follow directly from Theorem 7.4 and [ES18, Theorem 1.4]. However, there is a little subtlety involved: the Green function constructed by Ehlen and Sankaran is a little different from that of Bruinier when $n=2$ as in our case. The next three subsections explains the subtlety and proves Theorem 1.1.
@@ 7.3. A little preparation
To define the Ehlen and Sankaran’s Green functions, we need a little preparation. In this subsection, let $L$ be an even dimensional integral non-degenerate $\mathbb{Z}$-lattice of signature $(2,2)$ (in our case, view our unimodular $\mathcal{O}_{\boldsymbol{k}}$-lattice $L$ of signature $(1,1)$ as a quadratic lattice with quadratic form $q(x)=(x,x)$). Let $L^{\prime}$ be the dual of $L$ with respect to the quadratic form, and $S_{L}=\mathbb{C}[L^{\prime}/L]$ with Weil representation $\rho_{L}$. Let $S_{L}^{\vee}$ be its dual with dual Weil representation $\rho_{L}^{\vee}$. Let $\phi_{\mu}$ ($\mu\in L^{\prime}/L$) be the standard basis of $S_{L}$ and $\phi_{\mu}^{\vee}$ be the dual basis of $S_{L}^{\vee}$.
@@
An element in $H_{k}(\rho_{L})$ is called a harmonic Maass modular form (of weight $k$ with values in $S_{L}$). It is called a weakly holomomorphic modular form if $\xi_{k}(F)=0$. We denote the space of weakly holomomorphic forms by $M_{k}^{!}(\rho_{L})$.
According to Bruinier and Funke ([BF04]). there is an exact sequence
(7.4) $0\rightarrow M_{k}^{!}(\rho_{L})\rightarrow H_{k}(\rho_{L})\rightarrow S_{2-k}(\rho_{L}^{\vee})\rightarrow 0$
@@
(7.6) $F^{+}(\tau)=\sum_{m>-\infty}c_{F}^{+}(m)q^{m},\quad c_{F}^{+}(m)\in S_{L}$
has only finitely many negative part, and its non-holomorphic part $F^{-}(\tau)$ is of exponential decay as $v=\operatorname{Im}(\tau)$ goes to infinity.
@@
[…]
@@
    where $\tilde{\phi}_{\mu}=\frac{1}{2}(\phi_{\mu}+\phi_{-\mu})$. We denote $c_{m}^{+}(0)=c_{m,0}^{+}(0,0)$ for the $\phi_{0}$-component of $c_{m,0}^{+}(0)$ for later use.
@@
    for every $f\in M_{0}(\rho_{L}^{\vee})$. Here $a_{f}(n,\nu)$ is the $(n,\nu)$-th coefficient of $f$, and $c_{m,\mu}^{+}(n,\nu)$ is the $(n,\nu)$-th coefficient of $F_{m,\mu}^{+}$.
@@
It follows from [ES18, Lemma 2.4] together with the following explicit construction of the splitting map $\eta$ in [ES18, (2.8)]. First notice that $M_{0}(\rho_{L})=S_{L}^{\operatorname{SL}_{2}(\mathbb{Z})}$ is the maximal subspace of $S_{L}$ on which $\operatorname{SL}_{2}(\mathbb{Z})$ acts trivially. Secondly, we have a natural embedding
@@
Let the notation be as in Lemma 7.7. Then $F_{m,\mu}=0$ for $m<0$, and $F_{m,\mu}\in M_{0}(\rho_{L})$.
Following Ehlen and Sankaran [ES18], we can view $F_{m}$ as a harmonic modular form of weight 00 valued in $S_{L}\otimes S_{L}^{\vee}$ via
@@
Let $V=L_{\mathbb{Q}}=L\otimes_{\mathbb{Z}}\mathbb{Q}$, and let $\mathcal{D}^{o}$ be the negative $2$-planes in $V_{\mathbb{R}}$. For $z\in\mathcal{D}^{o}$, decompose
@@
(7.8) $\theta(\tau,z,h;\phi)=v\sum_{x\in V}\phi(h^{-1}x)\phi_{\infty}(\sqrt{v},x,z)$
which is a modular form of $\tau$ of weight 00, and $\mathrm{SO}(V)(\mathbb{Q})$-invariant. Here
@@
which is a (non-holomorphic) modular form of $\operatorname{SL}_{2}(\mathbb{Z})$ of weight 00 valued in $S_{L}^{\vee}$. The Bruinier Green functions (the Ehlen and Sankaran version) are defined as
[…]
@@
is the truncated fundamental domain for $\operatorname{SL}_{2}(\mathbb{Z})$, and finally $\operatorname{CT}_{s=1}(f(s))$ is the constant term of the Laurent series of $f(s)$ at $s=0$.

### v2

@@ 6.2. Arithmetic special divisors
We regard a Green function as a metric on the corresponding line bundle via (6.3) and reduce the improper pullback to a proper pullback following an idea in [BHY15, §7.4]. Let $\widehat{\mathcal{L}}_{m}^{\diamond,\mathrm{tot}}$ be the metrized line bundle associated to $\widehat{\mathcal{Z}}^{\diamond,\mathrm{tot}}(m,v)$ with canonical section $\mathbf{1}_{m}^{\diamond}$ such that
$\log\|\mathbf{1}_{m}^{\diamond}(z)\|^{2}=G^{\diamond}(m,v)(z)$
for all $z\in\mathcal{D}^{\diamond}-Z^{\mathrm{tot}}(m,v)$. We use the notations $\widehat{\mathcal{L}}_{m}^{\mathrm{tot}}$ and $\mathbf{1}_{m}$ similarly. Recall that $\widehat{\Omega}^{\diamond}$ (resp. $\widehat{\Omega}$) is equipped with the natural metric such that
$\|s_{z}\|^{2}=-4\pi ve^{\gamma}(s_{z},s_{z}),$
and
(6.14) $\varphi_{\Lambda}^{*}(\widehat{\Omega}^{\diamond})=\widehat{\Omega}$.
Here by $\varphi_{\Lambda}^{*}$ we mean pullback as metrized line bundle, which is always defined. Finally, let $\widehat{\mathcal{L}}_{0}^{B}$ be the metrized line bundle associated to $(\mathcal{B}(0,v),\mathrm{Gr}(0,v))$ with a canonical section $\mathbf{1}_{0}^{B}$ such that
$\log\|\mathbf{1}_{0}^{B}(z)\|=G(0,v)(z).$
Now by (6.14), the desired identity is equivalent to
(6.15) $\varphi_{\Lambda}^{*}(\widehat{\mathcal{L}}_{m}^{\diamond,\mathrm{tot}}\otimes(\widehat{\Omega}^{\diamond})^{\otimes r_{\Lambda}(m)})=(\widehat{\mathcal{L}}_{0}^{B})^{\otimes r_{\Lambda}(m)}\otimes\bigotimes_{\begin{subarray}{c}m_{1}+m_{2}=m,\\ m_{1}\neq 0\end{subarray}}(\widehat{\mathcal{L}}_{m_{1}}^{\mathrm{tot}})^{\otimes r_{\Lambda}(m_{2})}.$
By Theorem 5.8, it holds at the line bundle level. So it suffices to prove that it holds for the metrics over $\mathcal{S}$ (all the metrics on $\mathcal{S}^{*}$ are natural extension of those on $\mathcal{S}$).
On the right hand side, the natural rational section
$s_{\mathrm{nice}}=(\mathbf{1}_{0}^{B})^{r_{\Lambda}(m)}\otimes\bigotimes_{\begin{subarray}{c}m_{1}+m_{2}=m,\\ m_{1}\neq 0\end{subarray}}\mathbf{1}_{m_{1}}^{\otimes r_{\Lambda}(m_{2})}$
has metric
(6.16) $\log\|s_{\mathrm{nice}}(y)\|^{2}=r_{\Lambda}(m)G(0,v)(y)+\sum_{\begin{subarray}{c}m_{1}+m_{2}=m,\\ m_{1}\neq 0\end{subarray}}r_{\Lambda}(m_{2})\mathrm{Gr}(m_{1},v)(y)$
for $y\in\mathcal{D}$.
On the left hand side, for $z\in\mathcal{D}^{\diamond}-\mathcal{D}$,
$-\log\|\mathbf{1}_{m}^{\diamond}(z)\|^{2}=\sum_{\begin{subarray}{c}0\neq x_{1}+x_{2}=x\in L^{\diamond}\\ x_{1}\in L,x_{2}\in\Lambda\\ Q(x)=m\end{subarray}}\mathrm{Gr}^{\diamond}(x,v)(z)$
$=\sum_{\begin{subarray}{c}0\neq x\in\Lambda\\ Q(x)=m\end{subarray}}\beta_{1}(4\pi vR^{\diamond}(x,z))+\sum_{\begin{subarray}{c}0\neq x_{1}\in L,Q(x_{1})=0\\ x_{2}\in\Lambda,\,Q(x_{2})=m\end{subarray}}\mathrm{Gr}^{\diamond}(x_{1}+x_{2},v)(z)+\sum_{\begin{subarray}{c}m_{1}+m_{2}=m\\ x_{1}\in L,\,Q(x_{1})=m_{1}\neq 0\\ x_{2}\in\Lambda,\,Q(x_{2})=m_{2}\end{subarray}}\mathrm{Gr}^{\diamond}(x_{1}+x_{2},v)(z).$
Recall that by (6.7) we have
$\beta_{1}(4\pi vR^{\diamond}(x,z))=-\gamma-\log(4\pi vR^{\diamond}(x,z))+f_{x}(z)$
where
$f_{x}(z)=-\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}\frac{(4\pi v\cdot R^{\diamond}(x,z))^{k}}{k!}.$
for all $z\in\mathcal{D}^{\diamond}$. Notice that $f_{x}(z)=0$ for $z\in\mathcal{D}$.
For $0\neq x\in\Lambda$, we have $(x,x)>0$ and we can define a local section $s_{x}$ of $\Omega^{-1}$ over $\mathcal{D}^{\diamond}$ so that over a point $z\in\mathcal{D}^{\diamond}$, $s_{x,z}$ is characterized by
$s_{x,z}(a)=(\mathrm{pr}_{z}(x),a)\quad\text{ for }a\in z.$
@@
(6.17) $\log\|s_{x,z}\|^{2}=-\gamma-\log(4\pi vR^{\diamond}(x,z)).$
So we have a rational section $s$ of $\Omega^{\otimes-r_{\Lambda}(m)}$ over $S$ with
$s(z)=\sum_{\begin{subarray}{c}0\neq x\in\Lambda\\ (x,x)=m\end{subarray}}s_{x,z}$
and
$\log\|s(z)\|^{2}=\sum_{\begin{subarray}{c}0\neq x\in\Lambda\\ (x,x)=m\end{subarray}}(-\gamma-\log(4\pi vR^{\diamond}(x,z)).$
As in [BHY15, §7.4], we have an analytic analogue of (5.23). More precisely, for $x\in\Lambda$, we have an isomorphism
(6.18) $\mathcal{O}_{Z(x)}|_{\mathcal{D}}\cong\Omega^{-1}$
so that the section $s_{x}$ of $\Omega^{-1}$ corresponds to the canonical section of $\mathcal{O}_{Z(x)}|_{\mathcal{D}}$. Moreover, by the same proof of [BHY15, Theorem 7.12], we have a section $s_{m}$ of $\Omega^{-1}$ on an étale neighborhood $U$ of $\mathcal{S}^{*}$ that corresponds to the canonical section of $\mathcal{O}_{\mathcal{Z}(m)}|_{U}$ under (5.23) and agrees with $s=\sum_{\begin{subarray}{c}0\neq x\in\Lambda\\ (x,x)=m\end{subarray}}s_{x}$ on $\mathcal{D}$. So we have a ‘proper’ rational section $s_{\mathrm{prop}}=\mathbf{1}_{m}^{\diamond}\otimes s^{-1}$ of $\mathcal{L}_{m}\otimes\Omega^{\otimes r_{\Lambda}(m)}$ that corresponds to $s_{\mathrm{nice}}$ under (6.15) and has metric
$-\log\|s_{\mathrm{prop}}(z)\|^{2}=-\log\|1_{m}^{\diamond}(z)\|^{2}+\log\|s(z)\|^{2}$
$=\sum_{\begin{subarray}{c}0\neq x\in\Lambda\\ Q(x)=m\end{subarray}}f_{x}(z)+\sum_{\begin{subarray}{c}0\neq x_{1}\in L,Q(x_{1})=0\\ x_{2}\in\Lambda,\,Q(x_{2})=m\end{subarray}}\mathrm{Gr}^{\diamond}(x_{1}+x_{2},v)(z)+\sum_{\begin{subarray}{c}m_{1}+m_{2}=m\\ x_{1}\in L,\,Q(x_{1})=m_{1}\neq 0\\ x_{2}\in\Lambda,\,Q(x_{2})=m_{2}\end{subarray}}\mathrm{Gr}^{\diamond}(x_{1}+x_{2},v)(z).$
Now for any $y\in\mathcal{D}$, let $z\in\mathcal{D}^{\diamond}-\mathcal{D}$ goes to $y$, the right hand side has limit, and since $f_{x}(z)=0$ for $z\in\mathcal{D}$, we have
$-\log\|s_{\mathrm{prop}}(y)\|^{2}$
$=\sum_{\begin{subarray}{c}0\neq x_{1}\in L,Q(x_{1})=0\\ x_{2}\in\Lambda,\,Q(x_{2})=m\end{subarray}}\mathrm{Gr}^{\diamond}(x_{1},v)(y)+\sum_{\begin{subarray}{c}m_{1}+m_{2}=m\\ x_{1}\in L,\,Q(x_{1})=m_{1}\neq 0\\ x_{2}\in\Lambda,\,Q(x_{2})=m_{2}\end{subarray}}\mathrm{Gr}^{\diamond}(x_{1},v)(y)$
$=-\log\|s_{\mathrm{nice}}(y)\|^{2}$
by (6.16) as desired. This proves the theorem.
@@ 7.1. Eisenstein Series
In this subsection, we temporarily let $L$ be a unimodular positive definite Hermitian lattice over $\mathcal{O}_{\boldsymbol{k}}$ of rank $\ell$. Let $\psi=\prod_{p}\psi_{p}$ be the canonical additive character of $\mathbb{Q}\backslash\mathbb{A}$ with $\psi_{\infty}(x)=e(x)=e^{2\pi ix}$. Associated to it is a Weil representation $\omega=\omega_{L,\psi}$ of $\hbox{SL}_{2}(\mathbb{A})\subset U(1,1)(\mathbb{A})$ on $S(V_{\mathbb{A}})$ with $V=L_{\mathbb{Q}}=L\otimes_{\mathbb{Z}}\mathbb{Q}$, and a $\operatorname{SL}_{2}(\mathbb{A})$-equivariant “Rallis” map ($s_{\ell}=\ell-1$)
$\lambda:\,S(V_{\mathbb{A}})\rightarrow I(s_{\ell},\chi),\quad\lambda(\phi)(g)=\omega(g)\phi(0).$
Here $\chi=\chi_{-D}^{\ell}$. Let $\phi_{L}=\mathrm{Char}(\widehat{L})\otimes\phi_{\infty}\in S(V_{\mathbb{A}})$ with $\phi_{\infty}(x)=e^{-\pi(x,x)}$, and let $\Phi=\Phi_{L}=\prod_{p\leq\infty}\Phi_{p}$ be the standard section in $I(s,\chi)$ such that $\Phi_{L}(g,s_{\ell})=\lambda(\phi_{L})(g)$. Then it is known that $\Phi_{\infty}=\gamma(V_{\infty})\Phi_{\infty}^{\ell}$ is the standard section of weight $\ell$ (up to a scalar multiplication by a local Weil index $\gamma(V_{\infty})$):
@@
Let (here $g_{\tau}=n(u)m(\sqrt{v})$ for $\tau=u+iv\in\mathbb{H}$)
(7.1) $E_{L}(\tau,s)=v^{-\ell/2}\sum_{\gamma\in B\backslash\operatorname{SL}_{2}(\mathbb{Q})}\Phi_{L}(\gamma g_{\tau},s),\quad$
be the associated Eisenstein series of weight $\ell$ (level $D$ and character $\chi$), which depends only on the genus $[[L]]$ of $L$. Here
@@
[…]
@@
$\Phi_{\infty}^{\ell}(\gamma g_{\tau},s_{\ell})=v^{\frac{\ell}{2}}j(\gamma,\tau)^{-\ell}.$
@@
    When $p\nmid D$, $\Phi_{p}(\gamma,s_{\ell})=1$ and $\Phi_{p}$ is right $\operatorname{SL}_{2}(\mathbb{Z}_{p})$-invariant.
@@
    $\Phi_{p}(\gamma,s_{\ell})=\begin{cases}\chi_{p}(a)&\hbox{if }p|c,\\ \gamma(V_{p})\chi_{p}(c)p^{-\ell/2}&\hbox{if }p\nmid c.\end{cases}$
@@
Notice that similar results hold when $s_{\ell}$ is replaced by general $s$ although we do not need it here. We have now the following theorem by the above lemmas.
@@
Let the notation be as above. Then for $\ell\geq 4$ and $s_{\ell}=\ell-1$,
[…]
@@ 7.2. Modularity of arithmetic theta functions with Kudla Green functions
Let the notation be as above. Then $\widehat{\Theta}_{K}(\tau)$ is a (non-holomorphic) modular form of $\Gamma_{0}(D)$ with values in $\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})$. By modularity, we mean that for every linear map $\lambda:\widehat{\operatorname{CH}}_{\mathbb{C}}^{1}(\mathcal{S}^{*})\rightarrow\mathbb{C}$, $\lambda(\widehat{\Theta}_{K}(\tau))$ is a real analytic modular form of the same type.
@@
By [BHK+20, Theorem B] and [ES18, Theorem 1.4], the left hand side of (7.3) is a (non-)holomorphic modular form of $\Gamma_{0}(D)$ of weight $2+\ell$ and character $\chi_{-D}^{\ell+2}$. So
$\widehat{\Theta}_{K}(\tau)=\frac{\sum_{\Lambda\in[[L_{0}^{\ell}]]}\frac{1}{|\mathrm{Aut}(\Lambda)|}\varphi_{\Lambda}^{*}(\widehat{\Theta}_{K}^{\diamond}(\tau))}{E_{L_{0}^{\ell}}(\tau,s_{\ell})}$
is a meromorphic modular form of $\Gamma_{0}(D)$ of weight $2$ and character $\chi_{-D}^{2}$ with possible poles at the zeros of the Eisenstein series. We now prove that $\widehat{\Theta}_{K}(\tau)$ is has no poles by varying the Eisenstein series. By Theorem 7.3, we see that
[…]
@@
Now, the modularity for $\widehat{\Theta}_{B}(\tau)$ (the associated Green functions are Bruinier Green functions as mentioned in the introduction) should follow directly from Theorem 7.4 and [ES18, Theorem 1.4]. However, there is a little subtlety involved: the Green function constructed by Ehlen and Sankaran is a little different from that of Bruinier when $n=2$ as in our case. The next three subsections explains the subtlety and proves Theorem 1.1.
@@ 7.3. A little preparation
We first recall Ehlen and Sankaran’s Green functions. In this subsection, let $L$ be an even dimensional integral non-degenerate $\mathbb{Z}$-lattice of signature $(2,2)$ (in our case, we view our unimodular $\mathcal{O}_{\boldsymbol{k}}$-lattice $L$ of signature $(1,1)$ as a quadratic lattice with quadratic form $q(x)=(x,x)$). Let $L^{\prime}$ be the dual of $L$ with respect to the quadratic form, and $S_{L}=\mathbb{C}[L^{\prime}/L]$ with Weil representation $\rho_{L}$. Let $S_{L}^{\vee}$ be its dual with dual Weil representation $\rho_{L}^{\vee}$. Let $\phi_{\mu}$ ($\mu\in L^{\prime}/L$) be the standard basis of $S_{L}$ and $\phi_{\mu}^{\vee}$ be the dual basis of $S_{L}^{\vee}$.
@@
An element in $H_{k}(\rho_{L})$ is called a harmonic Maass modular form (of weight $k$ with values in $S_{L}$). It is called a weakly holomomorphic modular form if $\xi_{k}(F)=0$. We denote the space of weakly holomomorphic forms in $H_{k}(\rho_{L})$ by $M_{k}^{!}(\rho_{L})$, and the space of modular forms in $H_{k}(\rho_{L})$ by $M_{k}(\rho_{L})$.
According to Bruinier and Funke ([BF04]), the $\xi$-operator gives an exact sequence
(7.4) $0\rightarrow M_{k}^{!}(\rho_{L})\rightarrow H_{k}(\rho_{L})\rightarrow S_{2-k}(\rho_{L}^{\vee})\rightarrow 0.$
@@
(7.6) $F^{+}(\tau)=\sum_{m\gg-\infty}c_{F}^{+}(m)q^{m},\quad c_{F}^{+}(m)\in S_{L}$
has only finitely many negative terms, and its non-holomorphic part $F^{-}(\tau)$ is of exponential decay as $v=\operatorname{Im}(\tau)$ goes to infinity.
@@
Let $F\in H_{0}(\rho_{L})$ with $c_{F}^{+}(m)=0$ for all $m<0$. Then $F\in M_{0}(\rho_{L})=S_{L}^{\hbox{SL}_{2}(\mathbb{Z})}$ is a holomorphic modular form of weight 0, i.e., an element in $S_{L}$ fixed by $\hbox{SL}_{2}(\mathbb{Z})$ via $\rho_{L}$.
@@
    where $\tilde{\phi}_{\mu}=\frac{1}{2}(\phi_{\mu}+\phi_{-\mu})$. We denote $c_{m}^{+}(0)$ for the $\phi_{0}$-component of $c_{m,0}^{+}(0)$ for later use.
@@
    holds for every $f\in M_{0}(\rho_{L}^{\vee})$. Here $a_{f}(n,\nu)$ is the $(n,\nu)$-th coefficient of $f$, and $c_{m,\mu}^{+}(n,\nu)$ is the $(n,\nu)$-th coefficient of $F_{m,\mu}^{+}$.
@@
The lemma follows from [ES18, Lemma 2.4] together with the following explicit construction of a splitting map $\eta$ in [ES18, (2.8)]. First notice that $M_{0}(\rho_{L})=S_{L}^{\operatorname{SL}_{2}(\mathbb{Z})}$ is the maximal subspace of $S_{L}$ on which $\operatorname{SL}_{2}(\mathbb{Z})$ acts trivially. Secondly, we have a natural embedding
@@
Let the notation be as in Lemma 7.7. Then $F_{m,\mu}=0$ for $m<0$, and $F_{0,\mu}\in M_{0}(\rho_{L})$.
Following Ehlen and Sankaran [ES18], we can view $F_{m}$ as a harmonic modular form of weight 0 valued in $S_{L}\otimes S_{L}^{\vee}$ via
@@
Let $V=L_{\mathbb{Q}}=L\otimes_{\mathbb{Z}}\mathbb{Q}$, and let $\mathcal{D}^{o}$ be the oriented negative $2$-planes in $V_{\mathbb{R}}$. For $z\in\mathcal{D}^{o}$, decompose
@@
(7.8) $\theta(\tau,z,h;\phi)=v\sum_{x\in V}\phi(h^{-1}x)\phi_{\infty}(\sqrt{v}x,z)$
which is a modular form of $\tau=u+iv$ of weight 0, and $\mathrm{SO}(V)(\mathbb{Q})$-invariant. Here
@@
which is a (non-holomorphic) modular form of $\operatorname{SL}_{2}(\mathbb{Z})$ of weight 0 valued in $S_{L}^{\vee}$. The Bruinier Green functions (the Ehlen and Sankaran version) are defined as
(7.10) $G_{ES}(m,\mu)$ $=\int^{reg}\langle F_{m,\mu},\theta_{L}\rangle d\mu(\tau)$
$=\operatorname{CT}_{s=0}\lim_{T\rightarrow\infty}\int_{\mathcal{F}_{T}}\langle F_{m,\mu},\theta_{L}\rangle v^{-s}d\mu(\tau).$
@@
is the truncated fundamental domain for $\operatorname{SL}_{2}(\mathbb{Z})$, and finally $\operatorname{CT}_{s=0}(f(s))$ is the constant term of the Laurent series of $f(s)$ at $s=0$.